### v1

# Learning to Deliberate: Meta-policy Collaboration for Agentic LLMs with Multi-agent Reinforcement Learning

Abstract
Multi-agent systems of large language models (LLMs) show promise for complex reasoning, but their effectiveness is often limited by fixed collaboration protocols. These frameworks typically focus on macro-level orchestration while overlooking agents’ internal deliberative capabilities. This critical meta-cognitive blindspot treats agents as passive executors unable to adapt their strategy based on internal cognitive states like uncertainty or confidence. We introduce the Meta-Policy Deliberation Framework (MPDF), where agents learn a decentralized policy over a set of high-level meta-cognitive actions: Persist, Refine, and Concede. To overcome the instability of traditional policy gradients in this setting, we develop SoftRankPO, a novel reinforcement learning algorithm. SoftRankPO stabilizes training by shaping advantages based on the rank of rewards mapped through smooth normal quantiles, making the learning process robust to reward variance. Experiments show that MPDF with SoftRankPO achieves a a 4–5% absolute gain in average accuracy across five mathematical and general reasoning benchmarks compared to six state-of-the-art heuristic and learning-based multi-agent reasoning algorithms. Our work presents a paradigm for learning adaptive, meta-cognitive policies for multi-agent LLM systems, shifting the focus from designing fixed protocols to learning dynamic, deliberative strategies.
Introduction
As the limitations of single-agent reasoning become apparent in complex tasks (Park et al. 2023; Hong et al. 2023), the research frontier is shifting toward multi-agent systems (MAS), where multiple large language model (LLM) agents engage in collaborative reasoning to improve accuracy and robustness (Wang et al. 2023; Srivatsa, Maurya, and Kochmar 2024). Their effectiveness often depends on well-designed collaboration protocols, such as multi-turn debates and peer review mechanisms (Du et al. 2023; Liu et al. 2024). However, static protocols prevent agents from adapting their behavior to the unique demands of a given problem and to evolving context (Du et al. 2024; Yue et al. 2025).
Figure 1: Evolving from fixed static debate to trainable meta-policy deliberation, with dynamic strategy adaptation for multi-agent collaboration.
Recent efforts to overcome the static protocol bottleneck have largely focused on dynamic, system-level optimizations (Zhao et al. 2023), which we categorize as macro-level orchestration. These methods optimize the connections between agents, but critically neglect the deliberative meta-cognition (Sun, Zhang, and Mathews 2006; Yeung and Summerfield 2012; Koriat 2015) within each agent. This Meta-Cognitive Blindspot prevents agents from differentiating between confident judgments and uncertain hypotheses, causing them to uniformly apply external protocols irrespective of their internal cognitive state, falling back to a Static Deliberation Strategy. That is, agents employ fixed or pre-programmed response behaviors, lacking a strategic faculty to adaptively choose among persisting, refining, or conceding actions based on context and internal confidence.
To address these deficiencies, we introduce the Meta-Policy Deliberation Framework (MPDF), which formulates multi-agent collaboration as a decentralized partially observable Markov decision process (Dec-POMDP). Unlike prior approaches relying on static heuristics or predefined rules, our framework equips each agent with a lightweight, decoupled policy network. Each agent explicitly reasons about its internal cognitive state and strategically selects from discrete deliberation acts, such as Persist, Refine, or Concede. This meta-cognitive capability enables agents to dynamically calibrate their behavior according to evolving confidence and situational context.
To robustly optimize this policy under sparse and noisy team-level feedback, we introduce SoftRankPO, a scale-resilient reinforcement learning algorithm. SoftRankPO stabilizes policy gradients by converting raw rewards into rank-based advantages derived from smooth Gaussian percentiles, effectively mitigating the reward-scale sensitivity and high-variance challenges. Our main contributions are:
  • A meta-policy framework formulating multi-agent LLM reasoning as a Dec-POMDP for adaptive coordination through learned decision-making dynamics.
  • SoftRankPO, a scale-resilient policy optimization algorithm that stabilizes learning via smooth rank-based advantage shaping, enabling low-variance and reliable convergence across diverse reward regimes.
  • Consistent accuracy gains over competitive baselines on six reasoning benchmarks by combining meta-cognitive coordination and scale-resilient optimization.
Related Works
Multi-Agent LLM Collaboration.
Recent advances in large language models (LLMs) have led to growing interest in multi-agent systems (Abdelnabi et al. 2023; Wu et al. 2024; Motwani et al. 2024; Ishibashi and Nishimura 2024; Trirat, Jeong, and Hwang 2024; Yan et al. 2025; Dai et al. 2025). Collaborative multi-agent LLM frameworks for reasoning and planning (Qian et al. 2023; Zhang et al. 2023; Zheng et al. 2023; Chen, Han, and Zhang 2024; Zhang et al. 2025a; Zu et al. 2025) mainly follow two paradigms. First, prestructured approaches rely on fixed coordination schemes (Liu et al. 2023), like debate (Du et al. 2023; Liu et al. 2024) and peer review pipelines. These systems adopt static topologies to guide communication (Du et al. 2023; Qian et al. 2024). Second, self-organizing methods adapt interactions during inference by restructuring collaboration graphs (Hu, Lu, and Clune 2024; Shang et al. 2024; Zhang et al. 2024a). Some frame workflow design as program synthesis over modules (Zhuge et al. 2024; Zhang et al. 2024b), or assign dynamic roles across agents (Hu et al. 2024; Yue et al. 2025). However, these methods mainly rely on heuristic control or static policies (Chan et al. 2023; Talebirad and Nadiri 2023; Wei, Shan, and Li 2025).
Multi-Agent Reinforcement Learning for LLMs.
Reinforcement learning (RL) for multi-agent systems aim to train coordinated, role-specialized agents beyond static prompting or supervised tuning (Chen et al. 2022a, b; Jia et al. 2025; Zhu et al. 2025; Peiyuan et al. 2024; Park et al. 2025; Liao et al. 2025; Han et al. 2025; Wan et al. 2025). These methods demonstrate that multi-agent RL can significantly improve collaboration and generalization over prompting-based systems (Jiang et al. 2025; Alsadat and Xu 2024; Lin et al. 2025). To support scalable optimization, frameworks like MARTI (Zhang et al. 2025b) and MHGPO (Chen et al. 2025a) adopt centralized coordination with decentralized policy learning. Others, such as ReSo (Zhou et al. 2025) and Verco (Li et al. 2024) introduce dynamic agent selection and interpretable communication protocols. Together, these works represent a growing trend toward training LLM-based MAS that learn to communicate and adapt under reinforcement objectives (Wang et al. 2024; Zhang et al. 2024c; Sun, Huang, and Pompili 2024; Chen et al. 2025b).
Figure 2: An overview of the Meta-Policy Deliberation Framework (MPDF). (a) Meta-Cognitive State Representation: Each agent constructs a cognitive state with decision schema, reasoning profile and introspective confidence. (b) Meta-Policy Deliberation: The agent’s meta-policy network selects one of three high-level actions: Persist (if confident), Refine (to improve its own solution), or Concede (to yield to a peer solution). (c) Environment Interaction: The chosen action leads to a new response, yielding a reward and transitioning the system to the next state. (d) SoftRankPO: a stable policy optimization RL algorithm.
Method
We formalize the multi-agent deliberation process as a decentralized partially observable Markov decision process (Dec-POMDP), defined by the tuple $\langle\mathcal{I},\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\Omega,\mathcal{O},\gamma\rangle$. Here, $\mathcal{I}$ is the set of $N$ agents, $\mathcal{S}$ is the unobserved global state space representing the true problem-solving state, and $\mathcal{A}=\times_{i\in\mathcal{I}}\mathcal{A}^{i}$ is the joint action space composed of individual agent actions from $\mathcal{A}^{i}$.
At each deliberation round $t\in\{0,\dots,T-1\}$, each agent $i\in\mathcal{I}$ receives a local observation $o_{t}^{i}\in\Omega$ and executes a deliberative action $a_{t}^{i}\in\mathcal{A}^{i}$ according to its policy $\pi^{i}_{\theta}(a_{t}^{i}|o_{t}^{i})$. The environment transitions based on the function $\mathcal{T}(s_{t+1}|s_{t},\mathbf{a}_{t})$, where $\mathbf{a}_{t}$ is the joint action. Subsequently, each agent receives a reward $r_{t}^{i}=\mathcal{R}^{i}(s_{t},\mathbf{a}_{t})$, shaped to reflect its marginal contribution to the team’s objective. The collective goal is to learn an optimal, decentralized policy $\pi^{*}=(\pi^{1*},\dots,\pi^{N*})$ that maximizes the expected discounted return:
$\pi^{*}=\arg\max_{\pi}\ \mathbb{E}_{\pi}\left[\sum_{t=0}^{T}\gamma^{t}\sum_{i\in\mathcal{I}}r_{t}^{i}\right]$ (1)
The primary research challenge, and the core of our contribution, lies in the specific instantiation of the observation space $\Omega$ and the action space $\mathcal{A}$. Traditional approaches often rely on high-dimensional, unstructured text, leading to sample inefficiency and a focus on superficial linguistic patterns. In contrast, we propose a novel Meta-Cognitive Deliberation Framework, detailed below, which defines these spaces at a higher level of strategic abstraction.
The Meta-Cognitive Deliberation Framework
Our framework is designed to overcome the Meta-Cognitive Blindspot and Static Deliberation Strategy shortcomings of existing multi-LLM-agent frameworks by equipping each agent with the ability to reason about cognitive state and to select from high-level strategic acts.
The Meta-Cognitive State Space
We define an agent’s local state representation $z_{t}^{i}$ on a structured, low-dimensional vector summarizing its meta-cognitive status. This design choice forces the policy to learn abstract strategic reasoning rather than superficial text matching. The state $z_{t}^{i}$ is composed of three key components, with full feature specifications and prompt templates detailed in Appendix B.
Decision Schema ($z_{ans}$) represents the agent’s conclusive output, providing a concise summary of what the agent believes. It is extracted directly from the agent’s textual response by parsing the final answer into a standardized format (e.g., a scalar value for mathematical problems or a one-hot vector for multiple-choice questions).
Reasoning Profile ($z_{prof}$) is a vector of self-reported features that summarizes how the agent reached its belief. These features are produced by the agent LLM in a structured format as part of its single generative step, capturing attributes like the number of reasoning steps, the logical operators employed (e.g., algebraic, arithmetic), and a self-assessed confidence score. This co-generation process yields meta-data without incurring additional inference cost.
Introspective Confidence ($z_{conf}$) is generated via a dedicated introspection step, where a critic model, specifically the same LLM backbone acting in a evaluation capacity, is prompted to assess the correctness of the agent’s full reasoning trace given the original problem. The critic’s categorical judgment (e.g., “Correct” or “Incorrect”) is then encoded into a vector, providing a signal for further deliberation.
An agent’s full observation $o_{t}^{i}$ for decision-making is composed of its own local meta-cognitive state and those of its peers: $o_{t}^{i}=(z_{t}^{i},\{z_{t}^{j}\}_{j\neq i})$. The policy network, as a core component of our deliberation framework, then utilizes a cross-attention mechanism to integrate its introspective assessment with the social context provided by its peers. Please see Appendix for the full policy module structure.
The Deliberative Action Space
To address the issue of static strategies, we move beyond simple communicative acts to a set of high-level, strategic deliberative acts, $\mathcal{A}^{i}$. These actions represent fundamental choices in a collaborative reasoning process, allowing the agent to learn a truly adaptive policy. The action space is defined as: Persist, deciding to maintain and defend one’s current solution against competing alternatives; Refine, an act of triggering an internal process of self-correction and iterative improvement; and Concede, a strategic deference to a peer’s solution based on an assessment of its higher quality.
By learning a policy $\pi^{i}_{\theta}(a_{t}^{i}|o_{t}^{i})$ over this action space, our agents develop a dynamic, context-dependent deliberation strategy. They learn not just how to solve the problem, but how to behave optimally as a member of a reasoning collective—deciding when to lead, when to follow, and when to rethink, unlocking more efficient and robust reasoning.
SoftRankPO: A Stable Policy Optimization Algorithm for Deliberation
Unlike in traditional control tasks with smooth feedback, the rewards derived from deliberation are inherently sparse, high-variance, and arise from heterogeneous action outcomes. The Refine action, in particular, which depends on the stochastic generative output of an LLM, introduces significant uncertainty into the learning process. This environment yields a reward signal that is often heavy-tailed and whose scale can vary by orders of magnitude across different problem instances, compromising standard policy gradient methods, including PPO and even the more advanced GRPO. The stability of these algorithms hinges on the quality of the advantage function.In practice, rewards across tasks vary by orders of magnitude, and even simple affine rescaling $R\mapsto\alpha R+\beta$ can lead to arbitrarily inflated or collapsed gradients. The gradient variance satisfies $\operatorname{Var}[\nabla\mathcal{L}]\propto\operatorname{Var}[R]$, making learning unstable under batch-specific reward shifts. Although techniques such as baseline subtraction or batch normalization attempt to control for variance, they do not eliminate the underlying dependence on reward scale, which remains a brittle foundation in collaborative systems where agents operate with heterogeneous criteria and subjective feedback.
SoftRank: Placing Order before Scale
We introduce SoftRankPO, a KL–regularised policy‑optimization method that aligns a trainable policy with ranked preferences. By divorcing learning signals from raw reward magnitudes, SoftRankPO avoids scale‑induced instability and eliminates the need for separate value critics, while still operating inside a familiar KL trust‑region.
Consider a state $s$ with reward vector $\mathbf{R}=(R_{1},\dots,R_{K})$ for the $K$ candidate actions. We first convert ordinal information into a standardised score through the mapping
$p_{i}=(\frac{\operatorname{rank}(R_{i})+\tfrac{1}{2}}{K})^{\tau},\qquad A_{i}=\frac{A_{i}^{\star}-\bar{A}^{\star}}{\mathrm{std}(\mathbf{A}^{\star})+\varepsilon},$ (2)
where $\Phi^{-1}$ denotes the inverse standard normal CDF of $\mathcal{N}(0,1)$, $A_{i}^{\star}=\Phi^{-1}(p_{i})$ and $\bar{A}^{\star}=\tfrac{1}{K}\sum_{j}A_{j}^{\star}$. The temperature parameter $\tau>0$ controls the contrast between high- and low-ranked actions, enabling smoother or sharper advantage shaping. Equation (2) leaves the relative ranking of rewards intact, removes any dependence on affine transformations of $\mathbf{R}$, and produces zero‑mean, low‑variance scores whenever $K$ is reasonably large.
Theorem 1 (Zero‑mean and bounded‑variance of SoftRank scores).
Let $\mathbf{R}=(R_{1},\dots,R_{K})$ be a reward vector with distinct entries, and let $A_{i}$ be the SoftRank scores obtained from Eq. (2) with temperature $\tau>0$. Then
[…]
Through this mechanism, we obtain a stable and reliable advantage signal. Next, we show how to incorporate this new signal into a constrained optimization objective.
KL‑constrained Rank-matching Objective
Let $\pi_{\mathrm{ref}}$ be a frozen reference policy obtained, for example, from supervised fine‑tuning. Under SoftRankPO, we maximize
$\begin{split}\max_{\pi_{\theta}}\;\mathbb{E}_{s\sim\mathcal{D}}\Bigl{[}\mathbf{A}(s)^{\!\top}\log\pi_{\theta}(\cdot\mid s)-&\beta\,\mathrm{KL}\bigl{(}\pi_{\theta}\;\|\;\pi_{\mathrm{ref}}\bigr{)}\\ &\quad\;+\;\lambda\,\mathcal{H}\bigl{(}\pi_{\theta}\bigr{)}\Bigr{]},\end{split}$ (3)
where $\beta>0$ controls the KL trust‑region and $\mathcal{H}$ denotes Shannon entropy. Rewriting in terms of the implicit KL reward, $R_{\theta}(s,a)=\beta\log\frac{\pi_{\theta}(a|s)}{\pi_{\mathrm{ref}}(a|s)}$. Exploiting $\sum_{i}A_{i}(s)=0$ shows that maximizing $J(\theta)$ is equivalent to minimizing the following rank-matching loss:
$\mathcal{L}_{\mathrm{SR}}(\theta)=\mathbb{E}_{s}\!\Bigl{[}\frac{1}{K}\sum_{i=1}^{K}\bigl{(}\,R_{\theta}(s,a_{i})-\bar{R}_{\theta}(s)-R(s,a_{i})+\bar{R}(s)\bigr{)}^{\!2}\Bigr{]}.$ (4)
Hence SoftRankPO learns logits whose relative shape mirrors that of the true ranked rewards, without requiring any partition function. SoftRankPO can be interpreted either as a KL‑regularised policy optimizer or as a rank‑regression procedure that aligns implicit rewards with true rewards.
Closed‑form Gradient and Structural Decomposition.
Unlike classical policy gradient methods that rely on stochastic returns or value estimators, SoftRankPO admits a fully differentiable loss function grounded in observable reward differences, enabling an explicit derivation of the gradient with respect to policy parameters while also revealing the intrinsic structure of the update rule.
Since the advantage $\mathbf{A}$ has zero mean, the partition function derivative vanishes and
$\nabla_{\theta}\mathcal{L}_{\mathrm{SR}}(\theta)=-\frac{1}{\beta}\sum_{s,i}\bigl{[}(R-\bar{R})-(R_{\theta}-\bar{R}_{\theta})\bigr{]}\,\nabla_{\theta}\log\pi_{\theta}(a_{i}|s).$ (5)
Equation (5) can be recognised as the gradient of the mean‑squared error in Eq (4). Rearranging terms yields the following triple decomposition:
$\mathcal{L}_{\mathrm{SR}}(\theta)$ $=-\mathbb{E}_{s,a}\!\bigl{[}(R-\bar{R})\log\pi_{\theta}\bigr{]}\;+\;\frac{1}{2}\,\mathbb{E}_{s}\!\bigl{[}\operatorname{Var}_{a}[\log\pi_{\theta}]\bigr{]}$
$\quad\;+\;\mathbb{E}_{s}\!\bigl{[}\operatorname{Cov}_{a}[\log\pi_{\theta},\log\pi_{\mathrm{ref}}]\bigr{]},$ (6)
which simultaneously captures advantage weighting, entropy‑variance control, and KL trust‑region coupling within a single self‑consistent loss.
Global Optimality under the KL Constraint.
The strict convexity of $\mathcal{L}_{\mathrm{SR}}$ in $R_{\theta}$ yields a unique solution.
Theorem 2 (Unique KL‑constrained optimum).
The loss $\mathcal{L}_{\mathrm{SR}}(\theta)$ admits the unique minimizer
$\pi_{\theta}^{\star}(a|s)=\frac{\pi_{\mathrm{ref}}(a|s)\exp\{\beta^{-1}R(s,a)\}}{\sum_{a^{\prime}}\pi_{\mathrm{ref}}(a^{\prime}|s)\exp\{\beta^{-1}R(s,a^{\prime})\}},$
satisfied by the KL‑constrained reward‑maximizing policy.
This result guarantees that SoftRankPO remains faithful to the same trust‑region geometry exploited by PPO and DPO, while enjoying robustness to reward scaling.
Variance control and convergence rate.
We compare single‑batch gradient variance with the Group‑Relative Policy Optimisation (GRPO) baseline and state a first‑order convergence rate under standard smoothness assumptions.
Theorem 3 (Variance Dominance and Convergence).
Let $\widehat{g}_{\mathrm{SR}}$ and $\widehat{g}_{\mathrm{GR}}$ denote stochastic gradients computed from the same mini‑batch for SoftRankPO and GRPO, respectively. If the centred reward variance satisfies $\sigma_{R}^{2}\geq 1$, then
$\operatorname{Var}\!\bigl{[}\widehat{g}_{\mathrm{SR}}\bigr{]}\leq\bigl{(}1+\tfrac{1}{K-1}\bigr{)}\,\operatorname{Var}\!\bigl{[}\widehat{g}_{\mathrm{GR}}\bigr{]}.$
Stochastic gradient descent with step sizes $\eta_{t}=\eta_{0}/\sqrt{t}$} achieves $\min_{0\leq t<T}\mathbb{E}\|\nabla_{\theta}\mathcal{L}_{\mathrm{SR}}(\theta_{t})\|^{2}=\mathcal{O}(T^{-1/2}).$
The algorithm remains efficient with $O(K\log K)$ time complexity; supporting proofs are detailed in Appendix A.
Consensus-Driven Differential Reward Shaping
To enable fine-grained credit assignment in decentralized deliberation, we design a reward framework grounded in the principle of marginal contribution. Instead of assigning scalar rewards directly based on outcome quality, we decompose the reward signal into two complementary components: local self-improvement and global influence on consensus.
The step-wise delta reward quantifies an agent’s immediate self-improvement, defined as the change in correctness of its own output before and after an action: $r_{\text{local}}=\mathbb{I}[\text{Correct}^{\text{after}}]-\mathbb{I}[\text{Correct}^{\text{before}}]$}, encouraging agents to increase their individual answer quality. The lookahead consensus reward captures an agent’s contribution to the final consensus, defined by its marginal impact on the outcome. Let $\text{Consensus}^{\text{full}}$ and $\text{Consensus}^{-i}$ denote the correctness of the final group decision with and without agent $i$’s input. The global reward is computed as $r_{\text{global}}=\mathbb{I}[\text{Consensus}^{\text{full}}]-\mathbb{I}[\text{Consensus}^{-i}]$}.
Together, the total reward $r=r_{\text{local}}+r_{\text{global}}$ provides structured, causally grounded supervision that aligns individual optimization with collective utility, enabling stable and targeted training of multi-agent deliberation policies.
Supervised Bootstrapping and MARL Fine-Tuning
Our training pipeline follows a two-stage process that integrates supervised guidance with variance-robust reinforcement learning, enabling stable convergence of the deliberation strategy. First, for robust initialization, we pre-train the policy network $\pi_{\theta}$ on an expert dataset $\mathcal{D}_{\text{sft}}=\{(s,a^{\star})\}$, where labels $a^{\star}$ are derived from oracle solutions, by minimizing a standard cross-entropy loss, yielding a strong reference policy $\pi_{\text{sft}}$. Subsequently, we generate a static offline corpus $\mathcal{D}_{\text{off}}$ by executing rollouts with this $\pi_{\text{sft}}$ policy. This corpus stores tuples of encountered meta-cognitive states and the resulting group-wise reward vectors, i.e., $\mathcal{D}_{\text{off}}=\{(s,\mathbf{R})\}$. Second, we fine-tune $\pi_{\text{sft}}$ on $\mathcal{D}_{\text{off}}$ using our SoftRankPO objective (Eq. 3), aligning the policy with the stable, rank-based advantages derived from the offline data. The KL-divergence term ensures that the policy remains within a trust region of the sensible SFT initialization.
Experiment
We evaluate our framework on three standard reasoning tasks for LLM agents. (i) Mathematical reasoning is assessed on GSM8K, MATH, AIME, and AMC. (ii) General reasoning is measured with the MMLU benchmark. (iii) Code generation is evaluated on HumanEval. For all data sets, we follow the original splits and report the official metrics: Solve Rate for GSM8K / MATH / AIME / AMC, Accuracy in MMLU and Pass@1 in HumanEval.
Baselines.
We compare MPDF against a comprehensive suite of baselines, including single-agent methods: Vanilla, CoT and SC; multi-agent debate frameworks: LLM-Debate, PHP and DyLan; and dynamic workflow systems: AgentPrune, AFlow and GPTSwarm.
Implementation Details.
Following common practice, we set the number of agents and rounds to 3 by default. We instantiate agents with Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, Qwen2.5-7B-Instruct and Qwen2.5-3B-Instruct. For all results, we report the average over 3 random seeds. Detailed hyperparameters and all settings are provided in Appendix C to ensure full reproducibility.
Model MA Type GSM8K MATH AIME AMC MMLU HumanEval Avg
Vanilla N Single 72.71 66.91 3.33 8.43 57.83 48.17 42.90
CoT N Single 74.00 68.18 3.33 12.05 61.60 51.83 45.17
SC N Single 80.21 73.09 4.44 12.05 68.64 56.10 49.09
PHP Y Debate 80.36 74.36 4.44 15.67 68.72 56.70 50.04
LLM-Debate Y Debate 83.47 78.91 5.56 18.07 67.41 58.54 51.99
DyLAN Y Debate 82.63 76.55 6.67 19.28 66.84 60.37 51.50
GPTSwarm Y Dynamic 84.07 73.64 5.56 15.66 69.48 57.32 51.14
AgentPrune Y Dynamic 84.53 79.45 4.44 14.46 68.82 57.92 51.60
AFlow Y Dynamic 83.40 75.63 3.33 10.84 69.07 62.16 50.74
Ours: SFT Y Deliberation 83.55 79.27 5.56 16.87 68.73 58.54 52.09
Ours: SFT + PPO Y Deliberation 82.79 79.82 4.44 13.25 68.15 56.62 50.85
Ours: SFT + GRPO Y Deliberation 85.52 81.45 7.78 20.48 70.22 60.98 54.40
Ours: SFT + SoftRankPO Y Deliberation 86.35 82.72 7.78 22.89 70.87 61.59 55.37
Table 1: Comparison of performance across reasoning tasks. “MA” denotes the use of multiple agents. “Avg” reports the average accuracy across all applicable benchmarks. Our method consistently outperforms both single-agent and multi-agent baselines.
Model LMA-8B LMA-3B Qwen-8B Qwen-3B
Vanilla 72.71 24.72 90.90 83.62
CoT 74.00 50.49 91.05 84.23
SC 80.21 54.51 93.33 88.17
PHP 80.36 62.47 92.87 86.81
LLM-Debate 83.47 76.19 93.56 87.34
DyLAN 82.63 76.65 93.25 87.79
GPTSwarm 84.07 68.84 92.49 86.50
AgentPrune 84.53 65.35 92.12 86.35
AFlow 83.40 59.67 90.22 82.11
Ours 86.35 80.52 94.47 89.77
Table 2: Performance on GSM8K across different LLM backbones. Results demonstrate consistent performance gains across model scales and pre-training settings.
Dataset LLM-Debate Ours Improvement (%)
GSM8K 8,608 6,154 28.53%
MATH 7,474 4,346 41.85%
AIME 19,321 14,077 27.15%
AMC 17,917 12,620 29.58%
MMLU 10,908 7,322 32.87%
HumanEval 3,054 2,616 14.34%
Total 67,282 47,135 29.96%
Table 3: Average total token count (input and output combined) per sample across multiple datasets.
Main Results
Our Multi-Agent Deliberation Framework Outperforms Baselines.
As shown in Table 1, our deliberation-based multi-agent framework achieves the highest average accuracy (55.37%) across six diverse reasoning benchmarks, outperforming both debate-style methods (e.g., DyLAN 51.50%) and dynamic approaches (e.g., AgentPrune 51.60%). Unlike single-agent methods (e.g., Vanilla, CoT, SC), which lack inter-agent coordination, and debate-style models, which suffer from unstable decision-making and weak convergence, our framework jointly optimizes decision-making and consensus formation to enhance multi-agent reasoning. While routing methods attempt to improve efficiency through static interaction graphs and heuristic specialization, they remain limited by their inability to support dynamic deliberation or mutual influence during execution. In contrast, our approach establishes a unified, meta-cognitive deliberation process that enables agents to collaborate iteratively toward shared goals, improving accuracy especially in high-complexity tasks such as AMC and AIME.
Figure 3: Accuracy curves during deliberative training on GSM8K (top) and MATH (bottom) using the LLaMA3-8B backbone. Our method consistently achieves stable and faster convergence across both benchmarks.
Consistent Gains Across Diverse LLM Backbones.
To verify that our framework generalizes beyond a single model family, we conduct additional experiments using LLaMA3 and Qwen2.5 as representative backbones. Table 2 summarizes the results, demonstrating consistent and robust accuracy gains across all settings. On LLaMA-3B, the weakest backbone in our evaluation, accuracy improves from 68.84% (GPTSwarm) to 80.52%, indicating that the learned meta-policy effectively compensates for limited model capacity. Scaling up to LMA-8B still yields a positive margin, lifting the best baseline from 84.53% to 86.35%. Switching to Qwen models is particularly informative, as these checkpoints already benefit from extensive instruction tuning; Qwen-8B alone reaches 91.05%, yet deliberative training boosts it further to 94.47%; and Qwen-3B accuracy increases from 84.23% to 89.77%.
Lower Token Cost with Competitive Effectiveness.
To facilitate comparison, we report results against the widely adopted LLM-Debate baseline. Our method achieves a substantial reduction in average token use (47.1k), measured as both input and output tokens, outperforming LLM-Debate (67.2k) while maintaining or surpassing its accuracy (Table 3). LLM-Debate operates by aggregating all agent responses and prompting the model to synthesize an answer in a single pass, but lacks a mechanism for adaptive filtering. As a result, it often generates redundant and unfocused interactions, inflating token cost without proportional gains in solution quality. In contrast, our method introduces meta-cognitive policies that guide agents to act reflectively. Through reinforcement learning over interaction trajectories, agents learn to retain high-confidence outputs and intervene only when revisions are likely to yield substantial gains. This behavioral shift is evident in the increased PERSIST action rate (78.8% vs. 19.1% in baseline SFT), as shown in Figure 7, a move toward confident retention.
Ablation and Analysis Studies
From Fine-Tuning to SoftRankPO: Multi-Agent RL for Deliberation.
Our framework transitions from supervised fine-tuning (SFT) to progressively more effective reinforcement learning (RL) strategies that enhance agent coordination. As shown in Table 1, even the SFT-only variant performs competitively (52.09%), highlighting the strength of deliberation-based reasoning. However, it lacks outcome-driven feedback, limiting its ability to refine decision-making. Vanilla PPO performs worse (50.85%) due to unstable credit assignment and poor convergence in our multi-agent setting. As shown in Figure 3, PPO struggles especially on harder tasks (e.g., -3.62% on AMC), where its value-based updates fail to provide stable guidance. GRPO mitigates this by using group-relative ranking, stabilizing updates through peer comparison and improving performance on tasks like GSM8K (+1.97%) and MATH (+2.18%), but its reliance on raw reward magnitudes leads to sensitivity under noisy or heavy-tailed feedback. With reinforcement signals aligned to the group’s preference order, SoftRankPO yields the best overall results (55.37%) and more stable convergence dynamics (Figure 3).
Communication Rounds and Number of Agents.
Figure 4 shows how communication rounds ($M$) and agent number ($N$) affect final performance. Increasing $M$ from 1 to 4 steadily improves accuracy, validating the importance of multi-turn deliberation. However, performance slightly drops at 5 rounds, suggesting diminishing returns and possible noise accumulation, highlighting the need to calibrate deliberation depth. As $N$ increases from 2 to 6, accuracy consistently improves. Unlike unstructured setups that suffer from conflict or stagnation, our reflective coordination continues to benefit from additional agents.
Figure 4: Our framework scales well with more rounds and agents, with diminishing returns beyond optimal settings.
Robustness to Reward Scale.
We conduct an ablation where the reward signal is rescaled by exponential factors $\{0.1,1.0,10.0\}$. As shown in Figure 5, GRPO suffers from noticeable instability: different scales induce diverging learning curves, and all variants exhibit a pronounced drop in performance after early gains. This behavior stems from GRPO’s reliance on perstate $z$-score advantages, where an outlier can dominate the local mean and variance, leading to unstable gradient updates. In contrast, SoftRankPO is robust across reward scales. Its accuracy consistently improves and stabilizes around the same level regardless of the reward scaling, with visibly smoother and more stable curves. This empirical observation aligns with our analysis in Theorem 3. By inducing a zero-mean, bounded-variance advantage shaped by standard Gaussian, SoftRankPO effectively decouples learning dynamics from the scale of rewards.
Figure 5: Stability under different exponential reward scales. SoftRankPO remains robustness across multiple scales.
Impact of Temperature Parameter $\tau$ in SoftRankPO.
We conduct a controlled study varying the temperature coefficient $\tau\in\{0.4,0.6,0.8,1.0,2.0,5.0\}$, which adjusts the contrast in rank-derived advantage signals. We observe stable convergence across all settings (Figure 6), confirming that SoftRankPO remains robust to a broad range of shaping intensities. Notably, setting $\tau=0.8$ yields the best accuracy (approximately 88%), improving upon the default $\tau=0.5$ (86.35%). Increasing $\tau$ sharpens the distribution of transformed ranks, thereby amplifying differences in advantage estimates and enhancing gradient informativeness. Theoretically, since $A_{i}^{\star}=\Phi^{-1}(p_{i}^{\tau})$, a larger $\tau$ spreads out the values of $p_{i}^{\tau}$, especially near the extremes, increasing the gradient magnitude of the CDF inverse and leading to stronger learning signals for high-ranked actions. However, excessively low values of $\tau$ compress the percentile spacing and may obscure meaningful preference distinctions. Thus, $\tau$ influences both learning stability and policy expressiveness by regulating the granularity of rank sensitivity.
Figure 6: Effect of parameter $\tau$ on SoftRankPO performance. The method remains stable across a broad range of values, while finer tuning yields improved accuracy.
Policy Shift: From Collaboration to Coordination.
As shown in Figure 7, SoftRankPO induces a significant behavioral shift: the PERSIST rate increases from 19.1% to 78.8%, while REFINE and CONCEDE jointly fall from 81% to 21%. This stark transition suggests that agents have learned to recognize and trust the high-quality outputs produced by the pretrained LLM. It reflects an emerging ability for self-assessment and confidence calibration. The resulting policy aligns with the principle of minimal intervention: instead of revising reflexively or adopting peer responses blindly, agents now act selectively, intervening only when they perceive a meaningful improvement. Such a pattern is directly encouraged by our reward shaping, which emphasizes marginal utility and penalizes redundant edits. Beyond accuracy, this leads to reduced token use (Table 3). Furthermore, a detailed case study is presented in Appendix C.
Figure 7: Policy shift toward coordination: agents increasingly favor PERSIST, reflecting stronger confidence and selective intervention driven by reward shaping.
Conclusion
In this paper, we present MPDF, a meta-policy deliberation framework for multi-agent LLM collaboration, enabling agents to strategically adjust their behavior based on cognitive states. Coupled with the scale-resilient SoftRankPO algorithm, our approach achieves robust policy learning and fosters efficient coordination. Extensive results validate the effectiveness of our approach. We observe a policy shift during MPDF learning from collaboration to coordinatoin, marking a transition from exploratory language cooperation under SFT to coordinated decision-making under RL.